OpenStudy (anonymous):

whats the solution and steps for 4x−8=5x−2(x+2)

OpenStudy (anonymous):

4x-8=5x-2(x+2) can be written as 4x-8=5x-2x-4. You distribute the -2 to the x and positive 2. -2(x) = -2x, -2(2) = -4. Then you solve for x

OpenStudy (anonymous):

4x-8=3x-4. (3x=5x-2x) +8 +8 4x=3x+4 -3x -3x x=4

OpenStudy (anonymous):

Thank you !